OpenStudy (anonymous):

find the limit as x approaches 4 from the right of: (will post equations)

OpenStudy (anonymous):

f(x)= 2x-9 if x<4

OpenStudy (anonymous):

(x^2)-3 if 4<=x<8 and

OpenStudy (anonymous):

4x+23 if x>=8

OpenStudy (zzr0ck3r):

well x does not approach 4 from the right on the first one because x<4 and x does not approach 4 from the right on the next one because x>=8 so only the second one does x approach from the right. so plug in some number very close to 4 from the right, like 4.00000000001. or just 4 will work here because there is no domain problems...

OpenStudy (zzr0ck3r):

4.0000000.....00000001^2-3 is pretty much the same thing as 4^2-3 = ?

OpenStudy (anonymous):

oh so 13....ok thanks. the 8 in the equations kinda confused me. Thank you!